### v1

# General Coded Computing in a Probabilistic Straggler Regime

Abstract
Coded computing has demonstrated promising results in addressing straggler resiliency in distributed computing systems. However, most coded computing schemes are designed for exact computation, requiring the number of responding servers to exceed a certain recovery threshold. Additionally, these schemes are tailored for highly structured functions. Recently, new coded computing schemes for general computing functions, where exact computation is replaced with approximate computation, have emerged. In these schemes, the availability of additional results corresponds to more accurate estimation of computational tasks. This flexibility introduces new questions that need to be addressed. This paper addresses the practically important scenario in the context of general coded computing, where each server may become a straggler with a probability $p$, independently from others. We theoretically analyze the approximation error of two existing general coded computing schemes: Berrut Approximate Coded Computing (BACC) and Learning Theoretic Coded Computing (LeTCC). Under the probabilistic straggler configuration, we demonstrate that the average approximation error for BACC and LeTCC converge to zero with the rate of at least $\mathcal{O}(\log^{3}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-3}})$ and $\mathcal{O}(\log^{4}_{\nicefrac{{1}}{{p}}}(N)\cdot{N^{-2}})$, respectively. This is perhaps surprising, as earlier results does not indicate a convergence when the number of stragglers scales with the total number of servers $N$. However, in this case, despite the average number of stragglers being $Np$, the independence of servers in becoming stragglers allows the approximation error to converge to zero. These theoretical results are validated through experiments on various computing functions, including deep neural networks.
I Introduction
Coded computing has emerged as a powerful tool to improve the reliability and security of distributed computing systems, particularly, in presence of slow servers, known as stragglers, which fail to deliver results within a reasonable time, and adversarial servers that maliciously or unintentionally alter their outputs arbitrarily [1, 2, 3, 4, 5, 6].
Initially, the primary objective of coded computing was to achieve exact recovery, often within the context of finite fields. Within this framework, researchers have developed highly efficient solutions, particularly for structured computations such as polynomials [5, 7, 8] and matrix multiplications [2, 9, 10, 11, 12]. These methods often impose a strict recovery threshold: if the number of non-straggler servers exceeds this threshold, exact recovery is achievable; otherwise, the computation process fails entirely [5, 9].
The limited scope of conventional coded computing, however, does not address the requirements of modern distributed machine learning applications. In ML applications, computation tasks often lack a specific structure. Additionally, computations are performed over real numbers, where approximate results are generally sufficient. This motivates the development of general coded computing, which can handle a wide range of computations and exploit the possibility of approximate recovery to reduce the required computing resources [6, 1]. Unlike exact recovery, approximate recovery does not rely on a strict recovery threshold. Instead, the approximation error decreases as the number of non-straggling servers increases, making this approach more practical for real-world scenarios where exact computation is often infeasible or unnecessary [6, 4, 1, 13]. In particular, in two schemes, namely Berrut Approximate Coded Computing (BACC[1] and Learning Theoretic Coded Computing (LeTCC[6], the approximation errors are respectively bounded by $\mathcal{O}(\frac{S^{4}}{N^{2}})$ and $\mathcal{O}(\frac{S^{3}}{N^{3}})$, where $N$ is the number of the servers and $S$ is number of stragglers.
In this context, a practical question arises: if each server has a probability $p$ of becoming a straggler, independent of the others, rather than imposing a limit of at most $S$ straggler servers, does the approximation error still converge to zero? If it does, what is the rate of convergence? One would naively argue that in this case, the number of stragglers is in average $pN$. Since $pN$ scales with $N$, the results of [1] and [6] does not guarantee any convergence of approximation error to zero. Therefore, maybe in this case, we do not have any convergence.
In this paper, we address this question and show that the above understanding is not correct. We theoretically analyze the approximation error in the presence of probabilistic stragglers for two general coded computing schemes of BACC and LeTCC. We demonstrate that the approximation errors of BACC and LeTCC schemes converge to zero at rates of at least $\mathcal{O}(\log^{3}_{\frac{1}{p}}(N)N^{-3})$ and $\mathcal{O}(\log^{4}_{\frac{1}{p}}(N)N^{-2})$, respectively. This result demonstrates that the independence of servers in becoming stragglers indeed contributes positively to the convergence of the error. In the end, we validate our theoretical findings through experiments on various computing functions, including deep neural networks (see Fig. 1).
The paper is organized as follows: Section II provides the preliminaries, and Section III defines the problem formulation. Section IV presents the main theoretical results, with proof sketches detailed in Section V. Lastly, Section VI discusses the experimental results.
Notations: In this paper, bold uppercase letters (e.g., $\mathbf{A}$) are used to denote matrices, while bold lowercase letters (e.g., $\mathbf{x}$) represent vectors. Coded matrices and vectors are indicated with a tilde, such as $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{A}}$. The set $\{1,2,\ldots,n\}$ is denoted by $[n]$, and the size of a set $S$ is written as $|S|$. The $i$-th component of a vector-valued function $\mathbf{f}$ is expressed as $f_{i}(\cdot)$. The first, second, and $k$-th order derivatives of a scalar function $f$ are represented as $f^{\prime}$, $f^{\prime\prime}$, and $f^{(k)}$, respectively. The $\ell_{2}$-norm of a vector $\mathbf{x}$ is denoted by $\left\lVert\mathbf{x}\right\rVert_{2}$, while the $L_{p}$-norm of a function $f(\cdot)$ over the interval $\Omega$ is written as $\left\lVert f\right\rVert_{{L}^{p}\left(\Omega\right)}$. The space $\mathcal{H}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ corresponds to the reproducing kernel Hilbert space (RKHS) of second-order Sobolev functions for vector-valued functions of dimension $m$ over $\Omega$. Additionally, $\mathcal{C}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ represents the space of all vector-valued functions of dimension $m$ that with a second derivative over the interval $\Omega$. For one dimension, we simply use $\mathcal{H}^{2}\left(\Omega\right)$ and $\mathcal{C}^{2}\left(\Omega\right)$.
II Preliminaries
II-A Coded Computing for Straggler Resistance
Consider a scenario with one master node and $N$ servers. The master node aims to use the servers to compute $\{\mathbf{f}(\mathbf{x}_{k})\}_{k=1}^{K}$ in a distributed manner, where $\mathbf{x}_{k}\in\mathbb{R}^{d}$. Here, $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ can be any function, ranging from a simple one-dimensional function to a sophisticated deep neural network, with $K,d,m\in\mathbb{N}$. However, some servers can potentially become stragglers, not finishing their tasks within the required deadline. Therefore, assigning each data point to a single server would not be an effective approach.
To address this issue, the master node uses coding. It sends $N$ coded data points to each server using an encoder function. Each coded data point is a combination of the original data points. Each server then applies $\mathbf{f}(\cdot)$ to its received coded data, returning these coded results to the master node. The objective for the master node is to use some decoder function to approximate $\mathbf{\hat{f}}(\mathbf{x}_{k})\approx\mathbf{f}(\mathbf{x}_{k})$, even when some servers act as stragglers. The redundancy built into the coded data and the corresponding coded results allows the master node to reconstruct the intended outputs, $\{\mathbf{f}(\mathbf{x}_{k})\}_{k=1}^{K}$, using the results of the non-stragglers.
More formally, a general coded computing framework can be described as a three-step process:
  1. 1.
    Encoding: The master node uses an encoder function $\mathbf{u}_{\textrm{enc}}:\mathbb{R}\to\mathbb{R}^{d}$ to map a set of fixed, distinct, and ordered points $\alpha_{1}<\alpha_{2}<\dots<\alpha_{K}\in\mathbb{R}$ (encoder mapping points) into data points $\{\mathbf{x}_{k}\}_{k=1}^{K}$ in $\mathbb{R}^{d}$. This mapping can be either approximate, i.e., $\mathbf{x}_{k}\approx\mathbf{u}_{\textrm{enc}}(\cdot)$ [6] or precise, i.e., $\mathbf{x}_{k}=\mathbf{u}_{\textrm{enc}}(\cdot)$ [1]. Then, the same encoder function $\mathbf{u}_{\textrm{enc}}(\cdot)$ is applied to another set of fixed, distinct, and ordered points $\beta_{1}<\beta_{2}<\dots<\beta_{N}\in\mathbb{R}$ (decoder mapping points). The master node subsequently sends the encoded data points $\tilde{\mathbf{x}}_{n}=\mathbf{u}_{\textrm{enc}}(\beta_{n})$ to each server $n$, where each $\tilde{\mathbf{x}}_{n}$ represents a combination of all initial points $\{\mathbf{x}_{k}\}_{k=1}^{K}$.
  2. 2.
    Computing: Each server $n\in[N]$ computes $\mathbf{f}(\tilde{\mathbf{x}}_{n})=\mathbf{f}(\mathbf{u}_{\textrm{enc}}(\beta_{n}))$ on their given input and sends the result back to the master node.
  3. 3.
    Decoding: The master node collects the available results $\{\mathbf{f}(\tilde{\mathbf{x}}_{v})\}_{v\in\mathcal{F}}$ from the set of non-straggler servers $\mathcal{F}$. It then employs a decoder function $\mathbf{u}_{\textrm{dec}}:\mathbb{R}\to\mathbb{R}^{m}$ to map the points $\{\beta_{v}\}_{v\in\mathcal{F}}$ to the servers’ results $\{\mathbf{f}(\tilde{\mathbf{x}}_{v})\}_{v\in\mathcal{F}}=\{\mathbf{f}(\mathbf{u}_{\textrm{enc}}(\beta_{v}))\}_{v\in\mathcal{F}}$. Again, the decoder mapping can be either approximate [6] or accurate [1]. Finally, the master node computes $\hat{\mathbf{f}}(\mathbf{x}_{k}):=\mathbf{u}_{\textrm{dec}}(\alpha_{k})$ as an estimate of $\mathbf{f}(\mathbf{x}_{k})$ for $k\in[K]$ leveraging the relationship that $\mathbf{u}_{\textrm{dec}}(\alpha_{k})\approx\mathbf{f}(\mathbf{u}_{\textrm{enc}}(\alpha_{k}))\approx\mathbf{f}(\mathbf{x}_{k})$.
II-B Coded Computing Frameworks for General Computing Function
The first coded computing framework designed for general high-dimensional computing functions was Berrut Approximation Coded Computing (BACC) introduced by [1]. In this work, Berrut’s rational interpolation [14] is employed for mapping during both the encoding and decoding stages. More specifically, the encoder and decoder functions are defined as follows:
$\mathbf{u}_{\textrm{enc}}(z)=\sum_{i=1}^{K}\frac{\frac{(-1)^{i}}{\left(z-\alpha_{i}\right)}}{\sum_{j=1}^{K}\frac{(-1)^{j}}{\left(z-\alpha_{j}\right)}}\mathbf{x}_{i},$ (1)
$\mathbf{u}_{\textrm{dec}}(z)=\sum_{v\in\mathcal{F}}\frac{\frac{(-1)^{v}}{\left(z-\beta_{v}\right)}}{\sum_{i\in\mathcal{F}}\frac{(-1)^{i}}{\left(z-\beta_{i}\right)}}\mathbf{f}\left(\mathbf{u}_{\textrm{enc}}\left(\beta_{v}\right)\right).$ (2)
Berrut interpolant benefits from high stability compared to other rational interpolations, as its denominator does not have poles for any set of interpolation points [14]. Additionally, both encoding and decoding mapping are precise, meaning that $\mathbf{u}_{\textrm{enc}}(\alpha_{k})=\mathbf{x}_{k}$ and $\mathbf{u}_{\textrm{dec}}(\beta_{v})=\mathbf{f}(\mathbf{u}_{\textrm{enc}}(\beta_{v}))=\mathbf{f}(\tilde{\mathbf{x}}_{v})$ for $k\in[K]$ and $v\in\mathcal{F}$.
The authors in [6] propose a new scheme called LeTCC, based on learning theory. They define an end-to-end loss function for the entire process and leverage learning theory to design the encoding and decoding mappings.
[…]
where parameters $\lambda_{\textrm{dec}}$ and $\lambda_{\textrm{enc}}$ are hyper-parameters that control the smoothness of the decoding and encoding functions, respectively.
III Problem Formulation
In contrast to previous problem settings in coded computing, which assume that at most $S$ servers are stragglers, this paper investigates a more practical scenario where each server has a probability $p$ of being a straggler.
Following the notation from [6], the estimator function for the coded computing scheme is denoted by $\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$. where, $\bm{\alpha}=[\alpha_{1},\dots,\alpha_{K}]^{T}$ and $\bm{\beta}=[\beta_{1},\dots,\beta_{N}]^{T}$ and $\mathcal{F}$ refers to the set of non-straggler servers.
We define a random variable $F_{N}$ that is distributed over the set of all subsets of $N$ servers. Consequently, the performance of a coded computing estimator $\mathbf{\hat{f}}(\mathbf{x}):=\mathbf{\hat{f}}_{\bm{\alpha},\bm{\beta}}[\mathbf{u}_{\textrm{enc}},\mathbf{u}_{\textrm{dec}},\mathcal{F}](\cdot)$, applied to the data points $\{\mathbf{x_{k}}\}_{k=1}^{K}$, can be evaluated using the following average approximation error:
$\mathcal{L}(\mathbf{\hat{f}})$ $:=\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{1}{K}\sum^{K}_{k=1}\left\lVert\mathbf{\hat{f}}(\mathbf{x}_{k})-\mathbf{f}(\mathbf{x}_{k})\right\rVert^{2}_{2}\right]=\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{1}{K}\sum^{K}_{k=1}\left\lVert\mathbf{u}_{\textrm{dec}}(\alpha_{k})-\mathbf{f}(\mathbf{x}_{k})\right\rVert^{2}_{2}\right].$ (5)
Our goal is to derive an upper bound for $\mathcal{L}(\mathbf{\hat{f}})$ and compare the performance of BACC and LeTCC schemes in probabilistic configuration of stragglers.
IV Main Results
In this section, we provide our theoretical guarantees on the performance of the LeTCC and BACC frameworks under a probabilistic straggler model, where each server becomes a straggler with probability $p$. We establish an upper bound and analyze the convergence rate for both frameworks. For simplicity, we focus on a one-dimensional computing function $f:\mathbb{R}\to\mathbb{R}$. The results for both schemes are extendable to higher dimensions.
Let us define two variables representing minimum and maximum consecutive distance of decoder mapping points.
$\Delta_{\textrm{max}}:=\underset{n\in\{0,\dots,N\}}{\max}\{\beta_{n+1}-\beta_{n}\},\ \Delta_{\textrm{min}}:=\underset{n\in[N-1]}{\min}\left\{\beta_{n+1}-\beta_{n}\right\},$ (6)
Without loss of generality, we assume that the points $\{\alpha_{k}\}_{k=1}^{K}$ and $\{\beta_{n}\}_{n=1}^{N}$ lie within the interval $\Omega:=(-1,1)$.
The following theorems provide an upper bound for the LeTCC and BACC schemes respectively, under the probabilistic straggler configuration.
Theorem 1.
Consider the LeTCC framework consisting of $N$ servers, where $u_{\textrm{enc}}(\cdot),u_{\textrm{dec}}(\cdot)\in\mathcal{H}^{2}\left(\Omega\right)$ and $\lambda_{\textrm{dec}}\leqslant N^{-4}$. Let $f(\cdot)$ be a $\nu$-Lipschitz function satisfying $|f^{\prime\prime}|\leq\eta$ for some $\eta>0$. Assume that each server becomes a straggler with probability $p:=1-q>0$. If $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$ for some constant $B>0$, then for any $\delta\in(0,1)$, there exist constants $C_{l}$ and $n_{0}$ such that for $N>n_{0}$, with probability at least $1-\delta$, the following holds:
$\mathcal{L}(\hat{f})\leqslant C_{l}\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{3}}{N^{3}}\cdot\left\lVert(f\circ u_{\textrm{enc}})^{\prime\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}+\frac{2\nu^{2}}{K}\sum_{k=1}^{K}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(x_{k})\right)^{2}.$ (7)
Theorem 2.
Consider the BACC framework consisting of $N$ servers with $f\in\mathcal{C}^{2}\left(\Omega\right)$. Assume that each server is a straggler with probability of $p:=1-q>0$. If there is a constant $B$ such that $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$, then for any $\delta\in(0,1)$, there exist constants $C_{b}$ and $n_{0}$ such that for $N>n_{0}$ with a probability of at least $1-\delta$, we have:
$\mathcal{L}(\hat{f}){\leqslant}C_{b}\cdot B\frac{\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)^{4}}{N^{2}}\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}.$ (8)
Corollary 1.
By applying Theorems 1 and 2, in the probabilistic straggler setting, the average approximation error converges to zero with high probability. The convergence rates are at least $\mathcal{O}(\frac{\log^{3}_{\frac{1}{p}}(N)}{N^{3}})$ for LeTCC and at least $\mathcal{O}(\frac{\log^{4}_{\frac{1}{p}}(N)}{N^{2}})$ for BACC.
Remark 1.
Note that, in the probabilistic setting, the average number of stragglers in $N$ servers is approximately $Np$. Therefore, on average, the number of stragglers is a fraction of $N$. The upper bounds for the average approximation error of the BACC and LeTCC schemes with at most $S$ stragglers, as proved in [6] and [1], are $\mathcal{O}(\frac{S^{3}}{N^{3}})$ and $\mathcal{O}(\frac{S^{4}}{N^{2}})$, respectively. These results suggest that the average approximation error does not converge when the number of stragglers is a fraction of $N$. However, Corollary 1 demonstrates that the probabilistic behavior of stragglers causes the scheme’s approximation error to converge to zero.
Some commonly used mapping points do not satisfy the conditions stated in Theorems 1 and 2. Notably, the first and second Chebyshev points—widely used for interpolation and adopted in coded computing frameworks [1]—are prominent examples. However, the following corollary demonstrates that a similar upper bound can still be established for Chebyshev points.
Corollary 2.
Consider the LeTCC framework with $\lambda_{\textrm{dec}}\leqslant N^{-6}$ and the BACC framework, both consisting of $N$ servers. Assume the decoder mapping points are the first and second Chebyshev points, respectively: $\{\alpha_{k}\}_{k=1}^{K}=\cos(\frac{(2k-1)\pi}{2K})$ and $\{\beta_{n}\}_{n=1}^{N}=\cos(\frac{(n-1)\pi}{N-1})$. Under a probabilistic straggler configuration, the same upper bounds for convergence rates as stated in Corollary 1 can still be achieved.
Figure 1: Log-log plot showing the error convergence rates of LeTCC and BACC for the functions $f(x)=x\sin(x)$ and the LeNet5 network, with $p=0.05$ and $p=0.1$ (top). The plot also includes a comparison between $p=0.1$ and the non-probabilistic configuration with $S=2$ (bottom).
V Proof Sketch
In this section, we provide a proof sketch for the main theorems, with the formal proofs available in [15]. We first bound the average approximation error for each of LeTCC and BACC schemes.
(I) Bounding the average approximation error of LeTCC: Using the decomposition technique introduced in [6], we add and subtract $f(u_{\textrm{enc}}(\alpha_{k}))$ and apply the triangle inequality to bound the approximation error:
$\mathcal{L}(\hat{f})\leqslant\underbrace{\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{2}{K}\sum^{K}_{k=1}\left(u_{\textrm{dec}}(\alpha_{k})-f(u_{\textrm{enc}}(\alpha_{k}))\right)^{2}\right]}_{\mathcal{L}_{\textrm{dec}}(\hat{f})}+\underbrace{\frac{2}{K}\sum^{K}_{k=1}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(\mathbf{x}_{k})\right)^{2}}_{\mathcal{L}_{\textrm{enc}}(\hat{f})}.$ (9)
Since $f(\cdot)$ is $\nu$-Lipschitz, we can derive an upper bound for $\mathcal{L}_{\textrm{enc}}(\hat{f})$:
$\mathcal{L}_{\textrm{enc}}(\hat{f})\leqslant\frac{2\nu^{2}}{K}\sum_{k=1}^{K}\left(f(u_{\textrm{enc}}(\alpha_{k}))-f(\mathbf{x}_{k})\right)^{2}.$ (10)
Next, we drive an upper bound for $\mathcal{L}_{\textrm{dec}}(\hat{f})$. Defining $h(z):=u_{\textrm{dec}}(z)-f(u_{\textrm{enc}}(z))$, we obtain $\mathcal{L}_{\textrm{dec}}(\hat{f})\leqslant 2\mathrm{E}_{\mathcal{F}\sim F_{N}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right].$ Given that $f(\cdot)$ has bounded first and second derivatives, we can prove that $h\in\mathcal{H}^{2}\left(\Omega\right)$. Applying Sobolev interpolation inequalities [16], we can derive an upper bound for $\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}$:
$\mathrm{E}_{\mathcal{F}\sim F_{N}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]\leqslant 2\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}}\cdot\mathrm{E}_{\mathcal{F}}\left[\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\right]^{\frac{1}{2}}.$
In the next step, we derive an upper bound for $\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}$ and $\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}$. Given the set of non-straggler servers $\mathcal{F}$, let us define two variables representing the maximum and minimum distances between consecutive mapping points of the decoder function within the non-straggler set: $\Delta^{\mathcal{F}}_{\textrm{max}}:=\underset{f\in\left\{0,\dots,|\mathcal{F}|\right\}}{\max}\left\{\beta_{i_{f+1}}-\beta_{i_{f}}\right\},\Delta^{\mathcal{F}}_{\textrm{min}}:=\underset{f\in\left\{1,\dots,|\mathcal{F}|-1\right\}}{\min}\left\{\beta_{i_{f+1}}-\beta_{i_{f}}\right\}.$ The solution to (4) is a spline function, known as a second-order smoothing spline [17, 18, 19]. The norm of the error and the norm of its first derivative are bounded as follows [20]:
$\left\lVert h\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}{\leqslant}H_{0}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot L,$ (11)
and
$\left\lVert h^{\prime}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\leqslant H_{1}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}\cdot L^{\frac{1}{2}}(1+\frac{L}{16})^{\frac{1}{2}},$ (12)
where $H_{0}$ and $H_{1}$ are constants. The term $L$ is given by $L=p_{2}\left(\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}\right)\cdot\frac{(N-|\mathcal{F}|)\Delta^{\mathcal{F}}_{\textrm{max}}}{4}\lambda_{\textrm{dec}}+D\cdot\left(\Delta^{\mathcal{F}}_{\textrm{max}}\right)^{4}$, where $D$ is a constant and $p_{2}(\cdot)$ is a second-order polynomial with constant and positive coefficients. If $\lambda_{\textrm{dec}}\leqslant N^{-4}$, we can show that $L\leqslant\mathcal{O}((\Delta^{\mathcal{F}}_{\textrm{max}})^{4})$. Therefore, there exists a constant $C>0$ such that:
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]\leqslant C\cdot D\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{4}\right]^{\frac{1}{2}}\cdot\sqrt{D}\mathrm{E}_{\mathcal{F}}\left[(\Delta^{\mathcal{F}}_{\textrm{max}})^{2}\right]^{\frac{1}{2}}\cdot\sqrt{H_{0}H_{1}}\left\lVert(f\circ u_{\textrm{enc}})^{{}^{\prime\prime}}\right\rVert^{2}_{{L}^{2}\left(\Omega\right)}.$ (13)
(II) Bounding the average approximation error of BACC:
In [21] it is shown that
$\left\lVert h\right\rVert_{{L}^{\infty}\left(\Omega\right)}\leqslant\Delta^{\mathcal{F}}_{\textrm{max}}(1+\mu)\cdot\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)},$
where $\mu:=\max_{1\leq i\leq|\mathcal{F}|-2}\min\left\{\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}$ and :=$\mathcal{F}:=\{\beta_{j_{1}},\dots,\beta_{j_{|\mathcal{F}|}}\}$. Applying Max-Min inequality, we have: $\mu\leqslant\min\left\{\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}\leqslant\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}.$ Thus:
$\mathrm{E}_{\mathcal{F}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right]\leqslant\mathrm{E}_{\mathcal{F}}\left[\left(\Delta^{\mathcal{F}}_{\textrm{max}}(1+\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}})\right)^{2}\right]\cdot\left(\sum^{2}_{i=1}\left\lVert(f\circ u_{\textrm{enc}})^{(i)}\right\rVert_{{L}^{\infty}\left(\Omega\right)}\right)^{2}.$ (14)
As a result, (13) and (14) show that the random variables $\Delta^{\mathcal{F}}_{\textrm{max}}$ and $\Delta^{\mathcal{F}}_{\textrm{min}}$ play a key role in determining the approximation error for both coded computing schemes. Let us define random variable $R_{\mathcal{F},N}$ as the maximum number of consecutive stragglers among the servers. Thus, we have:
$\Delta^{\mathcal{F}}_{\textrm{max}}\leqslant(R_{\mathcal{F},N}+1)\cdot\Delta_{\textrm{max}},\quad\Delta^{\mathcal{F}}_{\textrm{min}}\geqslant\Delta_{\textrm{min}}.$ (15)
Therefore, establishing an upper bound for $\Delta^{\mathcal{F}}_{\textrm{max}}$ and $\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}$ is directly related to deriving an upper bound for $R_{\mathcal{F},N}$. Let us define i.i.d. Bernoulli random variables ${B_{1},\dots,B_{N}}$, where $B_{i}=1$ indicates that the server $i$ is straggler. Thus, $\mathrm{E}[B_{i}]=p$ for $i\in[N]$. Consequently, $R_{\mathcal{F},N}$ represents the maximum length of consecutive ones in the sequence ${B_{1},\dots,B_{N}}$. This random variable is commonly referred to as the longest head run in the literature [22, 23, 24, 25, 26].
Lemma 1.
([26, Theorem 2]) Let $R_{\mathcal{F},N}$ be the maximal sub-sequence length of consecutive ones or length of longest run in an i.i.d. Bernoulli sequence $B_{1},\dots,B_{N}$ with $\mathrm{E}[B_{i}]=p=1-q$ for $i\in[N]$. Let $\theta=\frac{\pi^{2}}{\ln\left(\frac{1}{p}\right)}$. Then:
$\left|\mathrm{E}\left[R_{\mathcal{F},N}\right]-\left(\log_{\frac{1}{p}}(qN)+\frac{\gamma}{\ln\left(\frac{1}{p}\right)}-\frac{1}{2}\right)\right|<g_{1}(\theta)+o(1),$
and
$\left|\operatorname{Var}\left[R_{\mathcal{F},N}\right]-\left(\frac{\pi^{2}}{6\left(\ln\left(\frac{1}{p}\right)\right)^{2}}+\frac{1}{12}\right)\right|<g_{2}(\theta)+o(1),$
where $\gamma$ is the Euler’s constant, $g_{1}(\theta):=\frac{\theta^{\frac{1}{2}}}{2\pi e^{\theta}(1-e^{-\theta})^{2}}$, and $g_{2}(\theta):=\frac{2(1.1+0.7\theta)\theta^{\frac{1}{2}}}{2\pi e^{\theta}(1-e^{-\theta})^{3}}$.
Lemma 1 states that although the $\mathrm{E}[R_{\mathcal{F},N}]$ grows with the $\log_{\frac{1}{p}}(N)$, its variance is nearly constant, i.e. $g_{2}(\theta)-(\frac{\pi^{2}}{6(\ln(\frac{1}{p}))^{2}}+\frac{1}{12})\leqslant\mathrm{Var[R_{\mathcal{F},N}]}\leqslant g_{2}(\theta)+(\frac{\pi^{2}}{6(\ln(\frac{1}{p}))^{2}}+\frac{1}{12})$. This suggests that $R_{\mathcal{F},N}$ is highly predictable, with its value approximately equal to $\log_{\frac{1}{p}}(qN)$. Leveraging this observation, it can be shown that there exist constants $n_{0}$ and $C$ such that for $N>n_{0}$, the following inequality holds:
$\Pr\left(R_{\mathcal{F},N}>C\left(\log_{\frac{1}{p}}(qN)+\sqrt{\frac{1}{\delta}}\right)\right)<\delta.$ (16)
By combining (16) with the results in (13) and (14), we can complete the proof.
VI Experimental Results
In this section, we evaluate the performance of the LeTCC and BACC schemes under the probabilistic straggler configuration. We consider two classes of computing functions for our experiments. The first is a one-dimensional function, $f_{1}(x)=x\sin(x)$. The second is a deep neural network as a high dimensional function. Specifically, we use the LeNet5 architecture [16] trained on a handwritten image classification task. The network takes a $32\times 32$ input image (flattened into a 1024-dimensional vector) and outputs a 10-dimensional vector representing the probability distribution over the classes. Thus, in this case, we are working with $\mathbf{f}_{2}:\mathbb{R}^{1024}\to\mathbb{R}^{10}$.
For the encoder and decoder mapping points, we use the first and second Chebyshev points, respectively as suggested by [1]. As an evaluation metric, we compute the empirical expectation of the recovery loss, $\mathrm{E}_{\mathbf{x}\sim\mathcal{X},\mathcal{F}\sim F_{N}}\left[\frac{1}{K}\sum^{K}_{k=1}\left\lVert\mathbf{u}_{\textrm{dec}}(\alpha_{k})-\mathbf{f}(\mathbf{x}_{k})\right\rVert^{2}_{2}\right]$. Specifically, for each value of $N$, we randomly sample input data points from the distribution and select stragglers with probability $p$. Each experiment is repeated $100$ times, and the average result is reported.
It is important to note that the smoothing parameters $\lambda_{\textrm{dec}}$ and $\lambda_{\textrm{enc}}$ do not impact the convergence rate in noiseless computations, as long as $\lambda_{\textrm{dec}}\leqslant N^{-4}$ (see [6]). Therefore, in our experiments, we use the non-smooth version of LeTCC, setting $\lambda_{\textrm{dec}}=\lambda_{\textrm{enc}}=0$.
Our experimental results are consistent with the theoretical findings presented in Section IV and reveal the following key observations: for both class of computing functions (1) the LeTCC scheme achieves a faster convergence rate compared to BACC, and (2) as $N$ increases, the convergence rate of average approximation error for the probabilistic configuration has a smaller exponent compared to the configuration with a maximum of $S$ stragglers (see Figure 1).
Acknowledgment
This material is based upon work supported by the National Science Foundation under Grant CIF-2348638.